Milk of magnesia is often taken to reduce the discomfort associated with heartburn. The recommended dose is one teaspoon, which contains 165 mg of Mg(OH)2. If the stomach contains 354 mL of pH 1.3 solution, what percent of the acid will be neutralized?
Answer #1

Mmol of H+ = M×V

= 0.05×354 =17.74 mmol

mmol of Mg(OH)2 = 165/58

= 2.85 mmol

mmol of OH- = 2×2.85 = 5.70 mmol

%Acid neutralize = (17.74-5.70)×100/17.74 = 68.0%